OpenStudy (anonymous):

Not too great with chemistry :/... Any help/guidance is greatly appreciated. 1. Observe the following reaction: NaOH + HCl → NaCl + H2O If 8.00 mL of HCl requires 7.52 mL of 0.10 M NaOH to consume the HCl (or titrate to completion), what is the concentration of HCl? 2. A 1 cm cuvette was filled with a solution which contains only a solute with a molar absorbtivity coefficient of 4800 M-1cm-1 and solvent. After blanking with the pure solvent the absorbance of the solution is 0.543. What is the concentration of the solution?

OpenStudy (anonymous):

1. lets write down all we have and all we need: NaOH + HCl → NaCl + H2O V(HCl)= 8 cm3 V(NaOH)= 7,52 cm3 c(NaOH)=0,10 M ------------------ c(HCl)=? so lets start: c=n/V -> n=c*V now we calculate amount of NaOH in moles: n(NaOH) = c(NaOH) * V(NaOH) n(NaOH) = 0,10 mol dm-3 * 7,52 * 10-3 dm3 n(NaOH) = ??? now you calculate what amount of NaOH is needed to completely react with HCl with bearing in mind ratio from balanced chemical equation: n(HCl) / n(NaOH) = 1/1 = 1 n(HCl) = n(NaOH) and now just calculate the concentration: c(HCl) = n(HCl) / V(HCl) any questions?